# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### 20 Feb 2024 v2

@@ Abstract
Quantum Error Correction Codes (QECCs) are fundamental to the advancement of quantum computing, safeguarding quantum states from the detrimental impact of noise and errors. The right choice of QECC, tailored to specific scenarios influenced by noise levels and qubit constraints, is as vital as the technological advancements themselves. This paper introduces a novel and comprehensive methodology for benchmarking QECCs, featuring a set of universal parameters. Utilizing eight distinguished QECCs, we propose a suite of eight parameters for a thorough analysis. Our work not only establishes a universal benchmarking methodology but also underscores the nuanced balance inherent in quantum error correction. The paper highlights that there is no one-size-fits-all solution; the selection of a QECC is contingent upon the specific constraints and circumstances of each case.
@@
Keywords: Quantum error correction codes (QECCs), Benchmarking, Parameters.
1 Introduction
@@
Quantum computing harnesses the principles of quantum mechanics to execute tasks that are beyond the capabilities of classical computing. This advanced technology finds applications in various fields, including molecule simulation for drug development, financial optimization enhancements, machine learning acceleration, improvement of optimization tasks, and revolutionary changes in supply chain management, among others [59, 55, 63, 31, 4]. However, commercialization of these quantum computing developments incurs challenges such as qubit stability and quantum noise [19, 48]. Quantum Error Correction Codes (QECCs) are crucial for achieving fault-tolerant quantum computing in the face of unavoidable qubit noise [46, 24]. Traditional error correction methods [34] encounter unique challenges in the quantum realm, primarily due to constraints like the no-cloning theorem [80] and the wave-function collapse that occurs during qubit measurement [77]. Research in this field has led to the creation of various quantum codes, such as the five-qubit code, Bacon-Shor code, topological code, surface code, color code, and heavy-hexagon code. These codes have shown effectiveness in correcting single errors [70, 2, 6, 41, 43, 11].
1.1 Motivation
Many quantum error-correcting codes have been proposed, each offering unique advantages and trade-offs. These range from the well-known Shor and Steane codes to more recent innovations like topological codes. However, this diversity leads to a critical question: “Which code is most suitable for a given quantum computing scenario?” This question underscores the need for a comprehensive, standardized benchmarking methodology. While numerous studies have individually assessed various QECCs, there is a lack of a unified methodology that allows for an objective, comparative analysis across codes. Such a comparison is vital not just for current quantum computing research but also for guiding future developments in QECCs.
1.2 Contribution
This paper introduces a comprehensive benchmarking methodology designed for all QECCs. Our methodology includes key parameters such as error correction capability, resource efficiency, error threshold, and scalability, ensuring a well-rounded evaluation of each code’s performance and practical feasibility. Through this methodology, we extensively compare various notable QECCs and their variations, offering insights into their behavior under different scenarios and highlighting their strengths and weaknesses. This methodology serves as a standard for assessing all QECCs and their variations, encompassing both existing and future ones.
To our knowledge, this is the first instance of a unified benchmarking methodology for all quantum error correction codes. With the ongoing evolution of quantum computing, new codes and their variations will continually arise. Our methodology offers a systematic approach to evaluate these codes, aiming to ensure that advancements in quantum computing are substantial and effectively address the quantum error challenge. In this study, we analyze eight prominent QECCs using eight distinct parameters, providing a clear comparative analysis. This study assists in choosing the most suitable QECC for their specific quantum computing applications.
1.3 Paper Structure
Section 2 begins with a concise overview of QECCs, followed by a detailed description of the eight QECCs utilized in this study. Section 3 introduces the eight benchmarking parameters, delves into the criteria for selecting codes & parameters, and outlines the methodology for comparative analysis. Section 4 presents an in-depth analysis of the eight QECCs in relation to the chosen parameters. Section 5 discusses prospective future work and challenges. Finally, the paper concludes in Section 6.
2 Theoretical Background
This section gives a general overview of the quantum error correction codes and then moves on to describe the codes that have been considered in this study.
2.1 Brief Overview of QECCs
2.1.1 Error Correction in Quantum Systems
Quantum Error-Correcting Codes (QECCs) differ fundamentally from classical error correction as they manage information encoded in quantum states, characterized by properties like superposition and entanglement [52]. Quantum systems are notably sensitive, with quantum states being easily disrupted by external disturbances [57]. This fragility presents a significant hurdle for reliable quantum computation and information storage, effectively addressed by QECCs [33].
QECCs function by encoding quantum information across multiple qubits, allowing the system to detect and correct errors without directly measuring the quantum state, in compliance with the no-cloning theorem [79, 25]. Quantum errors primarily include bit-flip and phase-flip errors, with more complex errors combining these two [65]. QECCs correct these errors through entangled states and collective measurements [13].
2.1.2 Code Varieties and Applications
Various QECCs have been proposed: Shor Code, the first QECC capable of correcting arbitrary single-qubit errors [65]; Steane Code, a Calderbank-Shor-Steane (CSS) code illustrating fault-tolerant quantum computation principles [67]; Topological Codes, like Kitaev’s Toric code, using a lattice of qubits for robustness against local errors [42] and Surface Codes, known for simpler error correction procedures, suitable for large-scale quantum computations [29].
QECCs are integral to practical quantum computing, particularly in achieving fault-tolerant quantum computation, a vital step for scalable and reliable quantum computers [3]. Ongoing research in QECCs focuses on optimizing code efficiency, reducing qubit overhead, and enhancing fault tolerance, crucial for the evolution of quantum computing [73].
2.2 Descriptions of QECCs Under Study
2.2.1 Repetition Code
The Quantum Repetition Code [24, 56] is one of the simplest forms of QECCs. Its primary function is to protect quantum information against errors, particularly bit-flip errors. This code extends the concept of the classical repetition code into the quantum realm, leveraging the fundamental principles of quantum mechanics [17]. The quantum repetition code works by encoding a single qubit of information across multiple qubits. For instance, in a 3-qubit repetition code, a single logical qubit or is encoded as or , respectively. This redundancy allows the code to detect and correct bit-flip errors that might occur on individual qubits.
@@ 2.2.1 Repetition Code
Error detection in the quantum repetition code is accomplished through a process known as syndrome measurement. This involves measuring the parity of neighboring qubits without collapsing the overall quantum state. For example, if the encoded state is and a bit-flip error occurs on the second qubit, the state changes to . The syndrome measurement detects this error by noticing the changed parity between the first and second qubits and the second and third qubits. Once detected, quantum operations can be applied to correct the error and restore the original state. Fig. 1 shows a quantum circuit for a 3-qubit repetition code, featuring state preparation, potential single bit-flip error detection, parity checking with two ancilla qubits, and error correction using self-inverse Pauli-gates.
2.2.2 Shor Code
The Shor Code [65], named after its inventor Peter Shor, is a landmark in the field of quantum error correction. Introduced in 1995, it was the first quantum error-correcting code (QECC) that demonstrated the feasibility of correcting arbitrary quantum errors, specifically both bit-flip and phase-flip errors, in a quantum bit or qubit.
The Shor Code encodes a single qubit of quantum information into a highly entangled state of nine physical qubits. This 9-qubit code is effectively a combination of the classical 3-bit repetition code, which corrects bit-flip errors, and the quantum phase-flip code, which corrects phase-flip errors.
@@ 2.2.2 Shor Code
The Shor Code utilizes a specific arrangement: The first layer of encoding replicates the qubit three times, preparing for bit-flip error correction and the second layer encodes each of these three qubits into three more qubits to guard against phase-flip errors. This two-layer structure makes the Shor Code capable of correcting any single-qubit error, whether it is a bit-flip, a phase-flip, or a combination of both on any of the nine qubits. It is described as a degenerate code, where different error types can affect the codewords similarly. The two basis states of the code are represented as and , intricately constructed from the superposition and entanglement of the nine qubits. The two basis states for the code are defined as: and . This structure ensures that the Shor Code can correct a single X error in any one block of three qubits and a single Z error on any of the nine qubits, classifying it as a full quantum error-correcting code
In the Shor Code, X errors are detected and corrected by using a correction circuit similar to that in the three-qubit code, applied to each block of three qubits. The process involves identifying single bit-flip errors and rectifying them accordingly. For phase-flip errors, the correction is achieved by examining the sign differences between the three blocks. This is executed using a set of six CNOT gates that compare the signs of blocks one and two, followed by another set of CNOT gates for blocks two and three. The unique aspect of the Shor Code is its degeneracy - a phase flip on any one qubit in a block of three has the same effect, hence, knowing the block in which the error occurs is sufficient to apply the correction operator. The Shor Code can correct for up to three individual bit flips provided each occurs in a different block of three. However, it is typically considered a single error-correcting code as it cannot handle multiple errors if they occur in specific locations. Fig. 2 depicts the encoding, error introduction, and error correction of a 9-qubit Shor code.
2.2.3 Steane Code
The Steane Code [66], also known as the 7-qubit code, was introduced by Andrew Steane in 1996 as a part of the Calderbank-Shor-Steane (CSS) [13, 67] family of quantum error-correcting codes. It represents a significant advancement in the field of quantum error correction, offering a symmetrical approach to correcting both bit-flip and phase-flip errors. The Steane Code encodes a single logical qubit into seven physical qubits. It is based on classical error-correcting codes, specifically the [7,4,3] Hamming code, which corrects single-bit errors in classical computing. In the quantum realm, this code is extended to correct both bit-flip and phase-flip errors in quantum bits. The key feature of the Steane Code is its CSS construction, which allows separate correction of bit-flip and phase-flip errors. This separation simplifies the error correction process and reduces the complexity of quantum circuits needed for implementation. The logical zero and one states in the Steane Code are defined using superpositions of the Hamming code’s even and odd weight codewords, respectively.
@@ 2.2.3 Steane Code
The Steane Code employs syndrome measurements for error detection, which uniquely indicate the presence and type of error without collapsing the quantum state. This code is comprised of 6 stabilizer generators: , , , , and . Due to this structure, the Steane Code can independently correct X and Z errors using separate circuits. This dual capability is central to its functionality, allowing for the efficient correction of single-qubit errors in quantum systems. The logical and states of this code are: and ; where . Fig. 3 depicts the error-detecting circuit of the 7-qubit Steane code.
2.2.4 Toric Code
The Toric Code [41], introduced by Alexei Kitaev, represents a novel approach to quantum error correction by leveraging topological properties. Its use of a two-dimensional lattice structure is unique, making it resistant to a broad class of local errors.
In the Toric Code, qubits are placed on the edges of a two-dimensional lattice on a torus. The topology of the lattice defines logical qubits and error correction is achieved by measuring vertex and plaquette operators. These operators are defined as: . In these expressions, and represent Pauli X and Z operators applied to the qubits. The products are taken over qubits around a vertex for vertex operators, and around a plaquette for plaquette operators.
@@ 2.2.4 Toric Code
Errors in the Toric Code manifest as excitations in this lattice structure. Bit-flip errors disrupt vertex operators, while phase-flip errors disrupt plaquette operators. The code detects errors by checking for these excitations. The topological nature of the code allows for the identification and correction of errors based on the collective state of multiple qubits, rather than individual qubit states. Fig. 4 represents the toric code and the acting of the X and Z stabilizers on its unique lattice structure.
2.2.5 Surface Code
Surface Codes [23] are a class of quantum error-correcting codes that extend the principles of the Toric Code to a planar geometry, meaning it is not modeled around a torus. They are known for their high threshold error rate, making them a promising candidate for scalable quantum computation.
@@ 2.2.5 Surface Code
Surface codes, similar to toric codes in stabilizers and error detection, have two main layouts: unrotated [29] and rotated [74]. Unrotated surface codes use a square lattice, with data qubits on square edges, Z-stabilizers on surfaces, and X-stabilizers at vertices. Rotated surface codes feature a tilted lattice, with X and Z stabilizers alternating in a chessboard pattern. The difference in lattice arrangement impacts their error-correction efficiency, with rotated codes generally preferred for long-distance QEC due to a higher error threshold and simpler implementation.
Error detection in Surface Codes is accomplished through the measurement of these stabilizer operators. A non-trivial measurement outcome (eigenvalue -1) indicates an error. The planar nature of Surface Codes facilitates the implementation of fault-tolerant gate operations and scalable quantum computing architectures. Fig. 5 shows the two different layouts of a surface code: unrotated and rotated.
2.2.6 Bacon-Shor Code
The Bacon-Shor Code [6, 65] is a type of quantum error-correcting code that simplifies error correction in quantum computing. It is a subclass of the more general family of Shor codes.
@@ 2.2.6 Bacon-Shor Code
The Bacon-Shor codes are stabilized over a square lattice, where the lattice dimensions dictate the X and Z error correction properties. The size of the lattice determines the total number of errors the code can correct. Such a code is defined as an lattice of qubits and is symmetric when . The X and Z stabilizers act on all qubits present on adjacent columns and rows respectively. If is an operator acting on the qubit at the position , which and . Tthe code stabilizer group is given by: , with generators expressed as a product of adjacent neighbor 2-qubit operators: and .
Syndrome extraction is done by measuring these operators, which are local and are on fewer qubits. The shortest Bacon-Shor code is with four stabilizer generators: , , and . Fig. 6 depicts the encoding and stabilizer measurement in Bacon-Shor code [69, 24].
2.2.7 3D Color Code
@@
The quantum color code [11], another form of topological QECC, takes a unique approach to error detection and correction. It extends the principles of surface codes by encoding logical qubits into a three-dimensional structure, often visualized as a two-dimensional lattice with ‘colors’ for ease of interpretation [44, 14]. There exist numerous two-dimensional renditions of a color code, among which we have chosen to utilize a representation in the form of a triangulated lattice filled with hexagonal surfaces for this discussion. Within this chosen format, each vertex signifies a logical qubit, while each hexagonal face corresponds to a stabilizer. However, color codes introduce a third type of stabilizer. In color codes, each hexagon in the lattice represents a different type of stabilizer: X-type, Z-type, and Y-type stabilizers. The Y-type stabilizer is associated with errors that involve both a bit-flip and a phase-flip simultaneously, thus accounting for its 3D structure. What sets color codes apart is their ability to detect and correct a broader range of errors. While surface codes can correct any single quantum error, theoretically color codes can handle more complex error models, including certain correlated or coherent errors. Fig. 7 showcases two distinct instances of color codes, each progressively larger as they scale according to distance. These color codes, while fundamentally three-dimensional, are depicted here within a two-dimensional triangular lattice format with hexagonal faces for ease of representation. The logical qubits composing the logical state are represented as gray-blobs. Each hexagonal face corresponds to a stabilizer, crucial to error detection and correction. Pink - hexagons denote Z-stabilizers, while the green - hexagons signify X-stabilizers. Differing from surface codes, color codes possess an additional stabilizer type, the Y-stabilizer, represented here by the yellow - hexagons.
3D Color Codes are capable of correcting a wider range of errors compared to their 2D counterparts. The introduction of an additional spatial dimension allows for more complex and robust error correction strategies. They can correct any arbitrary single-qubit error within their lattice structure and are particularly adept at dealing with correlated errors, which are a significant challenge in quantum computing.
Fig. 7 illustrates two iterations of color codes within a two-dimensional triangular lattice, simplifying their inherent three-dimensional structure. The diagram features logical qubits as gray blobs, surrounded by hexagonal stabilizers crucial for error detection and correction: pink hexagons for Z-stabilizers, green for X-stabilizers, and uniquely for color codes, yellow hexagons representing Y-stabilizers. This visual representation highlights the complex interplay and scalability of stabilizers in color codes, crucial for advanced quantum error correction.
2.2.8 Heavy Hexagon Code
Heavy hexagon codes [15] are a type of quantum error-correcting code that utilizes a lattice structure resembling a hexagon. They are designed to protect quantum information against errors more efficiently than some other quantum codes due to their unique geometric arrangement of qubits.
The heavy hexagon lattice is constructed by augmenting a hexagonal lattice with additional qubits placed at strategic locations. These additional ‘heavy’ qubits increase the code’s tolerance to errors by enhancing the connectivity between qubits, which is essential for the syndrome measurement process in error correction. The code’s error correction mechanism is designed to detect and correct both bit-flip (X) and phase-flip (Z) errors. It does so through a series of syndrome measurements that diagnose the type and location of the errors, allowing for targeted correction that preserves the fidelity of the quantum state.
This lattice stabilizer subsystem code synthesizes elements from both Bacon-Shor and surface-code frameworks, and encodes a single logical qubit into physical qubits where denotes the code distance. This lattice is characterized by its sparse connectivity, typically not exceeding three connections per qubit, making it a robust candidate for systems with fixed-frequency transmon qubits that are vulnerable to frequency collision errors.

### 05 Mar 2024 v3

@@ Abstract
Quantum Error Correction Codes (QECCs) are pivotal in advancing quantum computing by protecting quantum states against the adverse effects of noise and errors. With a variety of QECCs developed, including new developments and modifications of existing ones, selecting an appropriate QECC tailored to specific conditions is crucial. Despite significant improvements in the field of QECCs, a unified methodology for evaluating them on a consistent basis has remained elusive. Addressing this gap, this paper presents the first benchmarking framework for QECCs, introducing a set of universal parameters. By evaluating eight prominent QECCs, we propose a comprehensive suite of eight parameters for their analysis. Our methodology establishes a universal benchmarking approach and highlights the complexity of quantum error correction, indicating that the choice of a QECC depends on the unique requirements and limitations of each scenario. Furthermore, we develop a systematic strategy for selecting QECCs that adapts to the specific requirements of a given scenario, facilitating a tailored approach to quantum error correction. Additionally, we introduce a novel QECC recommendation tool that assesses the characteristics of a given scenario provided by the user, subsequently recommending a spectrum of QECCs from most to least suitable, along with the maximum achievable distance for each code. This tool is designed to be adaptable, allowing for the inclusion of new QECCs and the modification of their parameters with minimal effort, ensuring its relevance in the evolving landscape of quantum computing.
Index Terms:
@@
Quantum error correction codes (QECCs), Benchmarking, Parameters.I Introduction
@@
Quantum computing harnesses the principles of quantum mechanics to execute tasks that are beyond the capabilities of classical computing. This advanced technology finds applications in various fields, including molecule simulation for drug development, financial optimization enhancements, machine learning acceleration, improvement of optimization tasks, and revolutionary changes in supply chain management, among others [1, 2, 3, 4, 5]. However, commercialization of these quantum computing developments incurs challenges such as qubit stability and quantum noise [6, 7]. Quantum Error Correction Codes (QECCs) are crucial for achieving fault-tolerant quantum computing in the face of unavoidable qubit noise [8, 9]. Traditional error correction methods [10] encounter unique challenges in the quantum realm, primarily due to constraints like the no-cloning theorem [11] and the wave-function collapse that occurs during qubit measurement [12]. Research in this field has led to the creation of various quantum codes, such as the five-qubit, Bacon-Shor, topological, surface, color, and heavy-hexagon code [13, 14, 15, 16, 17, 18].
I-A Motivation
The landscape of quantum computing is rich with an array of QECCs, each heralding its unique set of advantages and compromises. From the foundational Shor and Steane codes to the cutting-edge realms of topological codes, the diversity of available QECCs presents a pivotal question: ”Which code is most suitable for a specific quantum computing scenario?” This query is at the heart of advancing quantum computing technologies and underscores the pressing need for a comprehensive and standardized benchmarking methodology. Despite the wealth of research assessing various QECCs, the absence of a unified framework for objective and comparative analysis across different codes stands as a significant gap. Such comparative insight is indispensable not only for the ongoing research within the quantum computing sphere but also for steering the direction of future QECC development.
Navigating through the myriad studies to deduce the optimal QECC for a particular scenario poses a formidable challenge, compounded further when multiple scenarios are considered. The complexity of determining the most appropriate QECC—factoring in the types of qubits, error rates, and other critical parameters—demands a systematic selection strategy. The necessity for a framework that can swiftly identify a suitable QECC based on specific scenario requirements is not just a matter of convenience but a crucial step forward in the practical implementation and advancement of quantum computing technologies.
I-B Contribution
To our knowledge, this research marks the inaugural effort to establish a unified benchmarking methodology for QECCs. Our objective is to discern how different QECCs are ideally suited for various scenarios and to devise a systematic strategy for the appropriate selection of QECCs. We aim to provide a comprehensive framework to facilitate this process. The contributions of this study are manifold and can be outlined as follows: Comprehensive Analysis of QECCs: We conduct an in-depth analysis of eight leading QECCs utilizing eight distinct parameters: qubit overhead, threshold, type of error protection, decoding complexity, types of transversal gates, scalability, the range of qubit types it has been realized on, and overall complexity. This enables a detailed comparative overview across different codes. Scenario-dependent QECC Suitability: Our findings emphasize that the selection of the most suitable QECC is contingent upon the specific scenario of its intended application, rather than a theoretical hierarchy of QECC efficacy. We demonstrate this through the use of radar charts, showcasing the trade-offs involved in choosing different QECCs, including variations within a single QECC type. Universal QECC Selection Strategy: We introduce a universal, systematic approach for selecting a QECC tailored to a given scenario. This strategy is designed to guide users in identifying the QECC that best matches their specific requirements, considering the multifaceted nature of quantum computing scenarios. An Adaptive QECC Recommendation Tool: A pivotal contribution of our work is the development of a first of its kind, dynamic tool that assesses user-provided scenario parameters and recommends a list of QECCs, ranked from most to least suitable, along with the maximum distance achievable for each code. This tool is designed to be evergreen, allowing for the inclusion of additional QECCs and the swift modification of their parameters without necessitating significant overhauls to the system. This ensures the framework’s adaptability and relevance in the rapidly evolving quantum computing landscape. For instance, should a QECC that has not been previously realized on trapped-ion qubits later achieve this milestone, the framework can easily accommodate this new information.
The tool, along with its code base, is available in a public GitHub repository. 111GitHub Repository Link: Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?
I-C Paper Structure
Section II provides a brief introduction to QECCs and an extensive examination of the eight QECCs featured in this research. Section III details the eight benchmarking parameters, selection criteria for codes and parameters, and the comparative analysis method. Section IV offers a thorough analysis of the eight QECCs against the selected parameters. Section V discusses a simplified code selection strategy and introduces our QECC recommendation tool. The paper concludes in Section VI.
II Theoretical Background
II-A Brief Overview of QECCs
QECCs differ fundamentally from classical error correction as they manage information encoded in quantum states, characterized by properties like superposition and entanglement [19]. Quantum systems are notably sensitive, with quantum states being easily disrupted by external disturbances [20]. This fragility presents a significant hurdle for reliable quantum computation and information storage, effectively addressed by QECCs [21]. QECCs function by encoding quantum information across multiple qubits, allowing the system to detect and correct errors without directly measuring the quantum state, in compliance with the no-cloning theorem [22, 23]. Quantum errors primarily include bit-flip and phase-flip errors, with more complex errors combining these two [24]. QECCs correct these errors through entangled states and collective measurements [25]. Various QECCs have been proposed: Shor Code, the first QECC capable of correcting arbitrary single-qubit errors [24]; Steane Code, a Calderbank-Shor-Steane (CSS) code illustrating fault-tolerant quantum computation principles [26]; Topological Codes, like Kitaev’s Toric code, using a lattice of qubits for robustness against local errors [27] and Surface Codes, known for simpler error correction procedures, suitable for large-scale quantum computations [28]. QECCs are integral to practical quantum computing, particularly in achieving fault-tolerant quantum computation, a vital step for scalable and reliable quantum computers [29]. Ongoing research in QECCs focuses on optimizing code efficiency, reducing qubit overhead, and enhancing fault tolerance, crucial for the evolution of quantum computing [30].
II-B Descriptions of QECCs Under Study
@@
Repetition Code: The Quantum Repetition Code [9, 31] is one of the simplest forms of QECCs. Its primary function is to protect quantum information against errors, particularly bit-flip errors. This code extends the concept of the classical repetition code into the quantum realm, leveraging the fundamental principles of quantum mechanics [32]. The quantum repetition code works by encoding a single qubit of information across multiple qubits. For instance, in a 3-qubit repetition code, a single logical qubit or is encoded as or , respectively. This redundancy allows the code to detect and correct bit-flip errors that might occur on individual qubits. Error detection in the quantum repetition code is accomplished through a process known as syndrome measurement. This involves measuring the parity of neighboring qubits without collapsing the overall quantum state. For example, if the encoded state is and a bit-flip error occurs on the second qubit, the state changes to . The syndrome measurement detects this error by noticing the changed parity between the first and second qubits and the second and third qubits. Once detected, quantum operations can be applied to correct the error and restore the original state. Fig. 1 shows a quantum circuit for a 3-qubit repetition code, featuring state preparation, potential single bit-flip error detection, parity checking with two ancilla qubits, and error correction using self-inverse Pauli-gates.
@@
Shor Code: The Shor Code [24], developed by Peter Shor in 1995, marks a significant advancement in quantum error correction by being the first to correct arbitrary quantum errors, including both bit-flip and phase-flip errors, in a qubit. It encodes one qubit of information into nine physical qubits through a two-layer structure: the first layer addresses bit-flip errors by tripling the qubit, and the second layer corrects phase-flip errors by further encoding each of these into three qubits. This innovative approach allows the Shor Code to correct any single-qubit error across its nine qubits. It is described as a degenerate code, where different error types can affect the codewords similarly. The two basis states of the code are represented as and , intricately constructed from the superposition and entanglement of the nine qubits. The two basis states for the code are defined as: and . This structure ensures that the Shor Code corrects one X error per trio of qubits and a single Z error across nine qubits, making it a comprehensive quantum error-correcting code. X errors are managed by a correction circuit for each trio, identifying and fixing bit-flips. Z errors are addressed by comparing sign differences between trios with CNOT gates. Its degeneracy allows it to correct a phase flip in any qubit of a trio effectively. Although capable of correcting three bit flips in separate trios, it’s primarily a single-error corrector due to its limitations with multiple errors. Fig. 2 shows the Shor Code’s encoding, error introduction, and correction process.
@@
Steane Code: The Steane Code [33], also known as the 7-qubit code, was introduced by Andrew Steane in 1996 as a part of the Calderbank-Shor-Steane (CSS) [25, 26] family of quantum error-correcting codes. It represents a significant advancement in the field of quantum error correction, offering a symmetrical approach to correcting both bit-flip and phase-flip errors. The Steane Code maps one logical qubit to seven physical ones, drawing from the [7,4,3] Hamming code to correct single-bit errors and extending it to fix both bit-flip and phase-flip errors in qubits. Its CSS construction enables independent correction of these errors, streamlining the correction process and simplifying the required quantum circuits. The logical zero and one states in the Steane Code are defined using superpositions of the Hamming code’s even and odd weight codewords, respectively. The Steane Code employs syndrome measurements for error detection, which uniquely indicate the presence and type of error without collapsing the quantum state. This code is comprised of 6 stabilizer generators: , , , , and . Due to this structure, the Steane Code can independently correct X and Z errors using separate circuits. This dual capability is central to its functionality, allowing for the efficient correction of single-qubit errors in quantum systems. The logical and states of this code are: and ; where . Fig. 3 depicts the error-detecting circuit of the 7-qubit Steane code.
@@
Toric Code: The Toric Code [16], introduced by Alexei Kitaev, represents a novel approach to quantum error correction by leveraging topological properties. Its use of a two-dimensional lattice structure is unique, making it resistant to a broad class of local errors. In the Toric Code, qubits are placed on the edges of a two-dimensional lattice on a torus. The topology of the lattice defines logical qubits and error correction is achieved by measuring vertex and plaquette operators. These operators are defined as: . In these expressions, and represent Pauli X and Z operators applied to the qubits. The products are taken over qubits around a vertex for vertex operators, and around a plaquette for plaquette operators. Errors in the Toric Code manifest as excitations in this lattice structure. Bit-flip errors disrupt vertex operators, while phase-flip errors disrupt plaquette operators. The code detects errors by checking for these excitations. The topological nature of the code allows for the identification and correction of errors based on the collective state of multiple qubits, rather than individual qubit states. Fig. 4 represents the toric code and the acting of the X and Z stabilizers on its unique lattice structure.
@@
Surface Code: Surface Codes [34] are a class of quantum error-correcting codes that extend the principles of the Toric Code to a planar geometry, meaning it is not modeled around a torus. They are known for their high threshold error rate, making them a promising candidate for scalable quantum computation. Surface codes, similar to toric codes in stabilizers and error detection, have two main layouts: unrotated [28] and rotated [35]. Unrotated surface codes use a square lattice, with data qubits on square edges, Z-stabilizers on surfaces, and X-stabilizers at vertices. Rotated surface codes feature a tilted lattice, with X and Z stabilizers alternating in a chessboard pattern. The difference in lattice arrangement impacts their error-correction efficiency, with rotated codes generally preferred for long-distance QEC due to a higher error threshold and simpler implementation. Error detection in Surface Codes is accomplished through the measurement of these stabilizer operators. A non-trivial measurement outcome (eigenvalue -1) indicates an error. The planar nature of Surface Codes facilitates the implementation of fault-tolerant gate operations and scalable quantum computing architectures. Fig. 5 shows the two different layouts of a surface code: unrotated and rotated.
@@
Bacon-Shor Code: The Bacon-Shor Code [15, 24] is a type of quantum error-correcting code that simplifies error correction in quantum computing. It is a subclass of the more general family of Shor codes. The Bacon-Shor codes are stabilized over a square lattice, where the lattice dimensions dictate the X and Z error correction properties. The size of the lattice determines the total number of errors the code can correct. Such a code is defined as an lattice of qubits and is symmetric when . The X and Z stabilizers act on all qubits present on adjacent columns and rows respectively. If is an operator acting on the qubit at the position , which and . Tthe code stabilizer group is given by: , with generators expressed as a product of adjacent neighbor 2-qubit operators: and . Syndrome extraction is done by measuring these operators, which are local and are on fewer qubits. The shortest Bacon-Shor code is with four stabilizer generators: , , and . Fig. 6 depicts the encoding and stabilizer measurement in Bacon-Shor code [36, 9].
@@
3D Color Code: The quantum color code [18], another form of topological QECC, takes a unique approach to error detection and correction. It extends the principles of surface codes by encoding logical qubits into a three-dimensional structure, often visualized as a two-dimensional lattice with ‘colors’ for ease of interpretation [37, 38]. In our discussion, we simplify color codes into a 2D triangular lattice with hexagons, where each vertex represents a logical qubit and each hexagon a stabilizer (X, Z, or the unique Y-type for simultaneous bit and phase flips). Color codes stand out for their ability to correct a wider array of errors, including complex correlated errors, beyond the single-error correction of surface codes. Fig. 7 illustrates this with two color code instances on a 2D lattice, showing logical qubits as gray blobs and stabilizers as hexagons: pink for Z, green for X, and yellow for Y, emphasizing the role of Y-stabilizers in their advanced error-correcting capabilities.
3D Color Codes offer superior error correction by leveraging an extra spatial dimension, enhancing their ability to fix single-qubit and correlated errors—a major quantum computing hurdle. Fig. 7 showcases color codes on a 2D triangular lattice, abstracting their 3D complexity. It depicts logical qubits (gray blobs) and hexagonal stabilizers: pink for Z, green for X, and yellow for Y stabilizers, illustrating the intricate stabilizer dynamics essential for effective error correction.